OpenStudy (anonymous):

The weight of 10,000 identical samples of a substance is 1 pound. What is the weight of 1000 samples?

OpenStudy (phi):

can you write down the ratio ?

OpenStudy (anonymous):

i dont know how to

OpenStudy (phi):

weight of 10,000 identical samples of a substance is 1 pound. write that as \[ \frac{ 10,000 \text{ samples} }{1 \text{ pound } }\]

OpenStudy (anonymous):

okay

OpenStudy (phi):

that ratio says 10,000 samples is "the same" as 1 pound the question asks for weight of 1000 samples \[ \frac{ 10,000 \text{ samples} }{1 \text{ pound } } = \frac{1000 \text{ samples} }{x \text{ pound } } \]

OpenStudy (phi):

that says 10,000 samples is to 1 pound as 1000 samples is to x (unknown) pounds to solve, "cross multiply" can you do that ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

1,000=x=10,000 ri ght ?

OpenStudy (phi):

there is only one = sign when you cross multiply try again.

OpenStudy (anonymous):

1,000=x10,000

OpenStudy (phi):

yes that is good.

OpenStudy (phi):

now we use the *very useful idea* anything divided by itself is 1 in other words divide both sides by 10,000 \[ \frac{1000}{10000} = \frac{x 10000}{10000} \\ \frac{1000}{10000} = \frac{x\cancel{ 10000}}{\cancel{10000}} \] we have \[ x = \frac{1000}{10000} \] can you simplify that ?

OpenStudy (anonymous):